OpenStudy (anonymous):

Write the equation of a line that passes through the point (-1,3) and is perpendicular to the line y=3

OpenStudy (anonymous):

what is the relationship with perpendicular lines and their slopes?

OpenStudy (anonymous):

the slopes are recipricol

OpenStudy (anonymous):

negative reciprocals.

OpenStudy (anonymous):

so you are given y = 3..the slope is ??

OpenStudy (anonymous):

-1/3

OpenStudy (anonymous):

y = mx + b slope is M

OpenStudy (anonymous):

so try it again, what is the slope of y = 3

OpenStudy (anonymous):

-1/3

OpenStudy (anonymous):

no the slope is 0.. think of this way - slopes are coefficients of the x variable when the y variable has a 1 in front of it.

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

so negative recip of 0 is 0 so with that information..use the point given (-1,3) to find the new equation y - y1 = m (x - x1) remember we just determined that M = 0

OpenStudy (anonymous):

so what is the new line?

OpenStudy (anonymous):

this is probably wrong but...y-3=x+1

OpenStudy (anonymous):

close :) you forgot the M

OpenStudy (anonymous):

y - 3 = 0 (x +1)

OpenStudy (anonymous):

and the final answer is where Y is on the side by itself..

OpenStudy (anonymous):

y=x-2?

OpenStudy (anonymous):

y=x+4

OpenStudy (anonymous):

actually, y = 3

OpenStudy (anonymous):

looks like you are not distributing properly...zero times a number is what?

OpenStudy (anonymous):

0

OpenStudy (anonymous):

exactly :)

OpenStudy (anonymous):

You good on this now?

OpenStudy (anonymous):

ok thanks again!

OpenStudy (anonymous):

No problem...I like helping you understand versus just wanting an answer. happy studying